OpenStudy (anonymous):

3x + y = 19 , and x + 3y = 1. Find the value of 2x + 2y ?

OpenStudy (kc_kennylau):

What do you get when you add those two equations together?

OpenStudy (kc_kennylau):

Divide the sum by 2 :D

OpenStudy (haseeb96):

3x+y=19......eq1 x+3y=1............eq2 adding them 4x+4y=20 2 common from both sides 2(2x+2y)=2(10) 2 cancel from both sides 2x+2y =10 this is the correct answer

OpenStudy (anonymous):

2x+2y=10 So that means 10 is the correct answer.